# Donsker–Varadhan large deviation principle for locally damped and randomly forced NLS equations

### v1

@@ 1.1. Main result
Let $\mathbf{B},\sigma>0$ be arbitrarily given. Assume the damping $a\colon\mathbb{T}\to[0,\infty)$ is smooth and not identically zero, and the noise $\eta$ can be formulated as in (2.5). Then there exists a constant $N\in\mathbb{N}^{+}$ such that if the non-negative numbers $b_{j,k}$ in (2.5) satisfy
$\sum_{j\in\mathbb{N}^{+},\,k\in\mathbb{Z}}|b_{j,k}|^{2}\langle k\rangle^{2(1+\sigma)}\leq\mathbf{B}.$ (1.2)
@@ 1.2. Literature review
The study of Donsker–Varadhan LDP for stochastic parabolic PDEs goes back to Gourcy [15, 16], based on earlier results for stochastic damped Hamiltonian system by Wu [37]. In these two paper, the LDP for the 1D viscous Burgers equation and 2D Navier–Stokes system was established, where the noise is white in time and sufficiently rough in space.
@@ 1.3. New ingredient
Here $X$ is a compact metric space, $\eta_{n}$ are random variables, and the map $S$ dictates the evolution. The LDP criterion relies on two assumptions, which can be roughly described as:
@@
We also mention that since $e^{-N^{2}t}$ decreases to 0 as $N$ tends to infinity, the squeezing rate can be arbitrarily small if $h$ is non-degenerate, by controlling sufficiently many modes using $h$ and let the sufficiently high frequencies decay by linear evolution.
@@ 1.4. Organization
$\mathcal{M}_{+}(X)$ and $\mathcal{P}(X)$ are the spaces of non-negative Borel measures and probability measures, both equipped with weak-* topology. $\mathcal{P}(X)$ is metricized by dual-Lipschitz distance:
@@ 2.2. Proof of the Main Theorem
To start with, let us state the precise setting on the noise structure. Denote by $\{\alpha_{j}:j\in\mathbb{N}^{+}\}$ an orthonormal basis of $L^{2}(0,1)$. We also define the trigonometric basis of $L^{2}(\mathbb{T})$ by
@@
Then $\{\alpha_{j}(t)e_{k}(k):j\in\mathbb{N}^{+},k\in\mathbb{Z}\}$ serves as an orthonormal basis of $L^{2}([0,1]\times\mathbb{T})$.
@@
    $\eta_{n}(t,x)=\eta(t+n,x)\quad\text{for }t\in[n,n+1),\ n\in\mathbb{N}_{0}$ (2.5)
    $\eta_{n}(t,x)=\chi(x)\sum_{j\in\mathbb{N}^{+},\,k\in\mathbb{Z}}b_{j,k}(\theta^{n}_{j,k,1}+i\theta_{j,k,2}^{n})\alpha^{\scriptscriptstyle T}_{j}(t)e_{k}(x),\quad t\in[0,T).$
    Here $\chi\in C^{\infty}(\mathbb{T})\setminus\{0\}$ is an arbitrary non-zero function333[7] assumed $\chi$ to be positive in an interval. But the proof still works out as long as $\chi$ is not identically zero., $b_{j,k}\geq 0$ are deterministic constants tending to 0 sufficiently fast (see (1.2)), and $\theta_{j,k,l}^{n}$ are independent real-valued random variables. Moreover, $\theta_{j,k,l}^{n}$ admits a probability density function $\rho_{j,k,l}$ supported by the interval [1,1][-1,1], which is $C^{1}$ and satisfies $\rho_{j,k}(0)>0$.
We point out that, under the boundedness condition (1.2), the support of noise $E:=\operatorname{supp}(\mathscr{D}(\eta_{n}))$ is a compact subset of $L^{2}(0,1;H^{1+\sigma}(\mathbb{T}))$, with respect to the the corresponding topology.
@@
$S\colon H^{1}\times L^{2}(0,T;H^{1})\to H^{1},\quad S(u,\zeta)=u(1).$
@@
Hypothesis $(\mathbf{C})$ has been demonstrated in [7, Section 5.3] via control theory, provided we equip $X$ with an equivalent norm (see [7, Lemma 4.10]). Meanwhile, a weaker version of $(\mathbf{I})$ is also derived in this prior work by global stability. Indeed, [7, Section 5.3] shown that (2.2) holds when $y=0$; that is, for any $\varepsilon_{0}>0$, one can find $N_{0}\in\mathbb{N}$ and $p_{0}>0$ such that
@@
and the sequence $\mathcal{A}_{n}$ is increasing. By compactness we see that $X\subset B_{X}(\mathcal{A}_{N_{1}}(\{0\}),\varepsilon_{1})$ for some $N_{1}\in\mathbb{N}$. In other words, for any $y\in X$, we can find $\zeta_{0},\dots,\zeta_{N_{1}-1}$ such that
@@
Thus, since $\zeta_{n}$ are i.i.d., we find that for any $z\in B_{X}(0,\delta)$,
@@
Since $\operatorname{supp}(\mathscr{D}(\eta_{n})=E\ni\zeta\mapsto\mathbb{P}(\eta_{n}\in B_{E}(\zeta,\delta))$ is lower-semicontinuous and pointwisely positive, and $E$ is compact, we conclude that $p_{1}$ is a positive constant. Finally, for any $\varepsilon>0$, let us set $\varepsilon_{1}=\varepsilon/2$, which determines the constant $\delta$, and then set $\varepsilon_{0}=\delta$. Then for $N:=N_{1}+N_{2}$ and any $x,y\in X$, by Markov property, we have
$P_{N}(x,B_{X}(y,\varepsilon))\geq P_{N_{1}}(x,B_{X}(0,\delta))\inf_{z\in B_{X}(0,\delta)}P_{N}(z,B_{X}(y,\varepsilon))\geq p_{0}p_{1}>0.$
@@ A $\Gamma$-type domain is a subdomain of $D$ in the form
    $\eta(t,x)=\eta_{n}(t-nT,x)\quad\text{for }t\in[n,(n+1)),\ n\in\mathbb{N}_{0},$ (2.7)
@@ 2.3.2. Navier–Stokes system
Finally, we investigate the 2D Navier–Stokes system with bounded kick force. In particular, our result extends [21, Theorem A] by allowing the noise to degenerate in high frequencies. The system in consideration are studied in, e.g., [26, 21, 22], which reads
@@
Here $D$ is a smooth bounded domain in $\mathbb{R}^{2}$, the constant $\nu>0$ is the viscosity, $u=(u^{1},u^{2})$ is the velocity field, and $p$ is the pressure. The phase space for (2.9) is specified as
@@
    $\eta(t,x)=\sum_{n=1}^{\infty}\delta(t-n)\eta_{n-1}(x),\quad\eta_{n}(x)=\sum_{k=1}^{\infty}b_{k}\theta_{k}^{n}\varphi_{k}(x)$ (2.10)
@@
Finally, we mention a similar application to Navier–Stokes system perturbed by degenerate bounded noise, which has been studied in [35] for exponential mixing. In this case the noise is akin to our setting for Schrödinger and wave equations. For the sake of simplicity, we omit the detailed description of this example, and claim that the LDP holds for the empirical distributions, as both $(\mathbf{I})$ and $(\mathbf{C}$) has essentially been verified in that work.
@@ 3.1. Some reductions
Reduction 1. Kifer’s criterion. In order to investigate the uniform LDP, let us consider the directed set [^6] A partially ordered set $(\Theta,\prec)$ is called an directed set, if any finite subset admits an upper bound. A net $\{x_{\theta}:\theta\in\Theta\}$ in a topological space refers to elements labeled by this directed set. The net $x_{\theta}$ converges to $x$, if for any neighborhood $U$ of $x$, there exists $\theta_{0}\in\Theta$ such that $x_{\theta}\in U$ whenever $\theta_{0}\prec\theta$. $\Theta:=\mathbb{N}\times\mathcal{P}(X)$, endowed with the partial order $\prec$ defined by
@@
which lower semicontinuous and convex. Moreover, $\Lambda$ can be reconstructed by the dual relation:
@@
Let $\Lambda\colon C(X)\to\mathbb{R}$ and $I\colon\mathcal{P}(X)\to[0,\infty]$ be defined as above. Given $V\in C(X)$, if the supreme on the right-hand side of (3.3) is attained at $\sigma_{V}\in\mathcal{P}(X)$, then $\sigma_{V}$ is called an equilibrium state for $V$.
@@
Assume the hypotheses $(\mathbf{I})$ and $(\mathbf{C})$ are satisfied, and $V\in L(X)$. Then there exists $\lambda_{V}>0$, $h_{V}\in C_{+}(X)$ and $\mu_{V}\in\mathcal{P}(X)$, such that for any $f\in C(X)$ and $\nu\in\mathcal{M}_{+}(X)$,
@@ 3.2. Lipschitz estimates for $Q_{n}^{V}$
Since $X$ is compact, so does $\mathcal{P}(X)$. Then Schauder’s fixed point theorem implies that, there exists a fixed point $\mu\in\mathcal{P}(X)$. In other words, let $\lambda_{2}=(Q^{V*}\mu)(X)>0$, we get (3.18).
@@ B.  Asymptotics of generalized Markov semigroups
Then there exists $\lambda>0$, $h\in C(X)$ and $\mu\in\mathcal{P}(X)$, such that for any $f\in C(Y)$ and $\nu\in\mathcal{M}_{+}(X)$,
@@
Acknowledgments   The authors would like to thank Ziyu Liu for valuable discussions and suggestions during the preparation of the paper. Yuxuan Chen would like to appreciate his PhD advisor Zhifei Zhang. Shengquan Xiang is partially supported by NSFC 12301562.

### v2

@@ 1.1. Main result
Let $\mathbf{B},\sigma>0$ be arbitrarily given. Assume the damping $a\colon\mathbb{T}\to[0,\infty)$ is smooth and not identically zero, and the noise $\eta$ can be formulated as in (2.5). Then there exists a constant $N\in\mathbb{N}$ such that if the non-negative numbers $b_{j,k}$ in (2.5) satisfy
$\sum_{j\in\mathbb{N},\,k\in\mathbb{Z}}|b_{j,k}|^{2}\langle k\rangle^{2(1+\sigma)}\leq\mathbf{B},$ (1.2)
@@ 1.2. Literature review
The study of Donsker–Varadhan LDP for stochastic parabolic PDEs goes back to Gourcy [15, 16], based on earlier results for stochastic damped Hamiltonian system by Wu [37]. In these two papers, the LDP for the 1D viscous Burgers equation and 2D Navier–Stokes system was established, where the noise is white in time and sufficiently rough in space.
@@ 1.3. New ingredient
Here $X$ is a compact metric space, $\eta_{n}$ are i.i.d. random variables, and the map $S$ dictates the evolution. The LDP criterion relies on two assumptions, which can be roughly described as:
@@
We also mention that since $e^{-N^{2}t}$ decreases to 0 as $N$ tends to infinity, the squeezing rate can be arbitrarily small if $h$ is non-degenerate, by controlling sufficiently many modes using $h$ and letting the sufficiently high frequencies decay by linear evolution.
@@ 1.4. Organization
$\mathcal{M}_{+}(X)$ and $\mathcal{P}(X)$ are the spaces of non-negative Borel measures and probability measures, both equipped with weak-* topology. $\mathcal{P}(X)$ is metrized by dual-Lipschitz distance:
@@ 2.2. Proof of the Main Theorem
To start with, let us state the precise setting on the noise structure. Denote by $\{\alpha_{j}:j\in\mathbb{N}\}$ an orthonormal basis of $L^{2}(0,1)$. We also define the trigonometric basis of $L^{2}(\mathbb{T})$ by
@@
Then $\{\alpha_{j}(t)e_{k}(x):j\in\mathbb{N},\ k\in\mathbb{Z}\}$ serves as an orthonormal basis of $L^{2}([0,1]\times\mathbb{T})$.
@@
    $\eta(t,x)=\eta_{n}(t-n,x)\quad\text{for }t\in[n,n+1),\ n\in\mathbb{N}_{0},$ (2.5)
    $\eta_{n}(t,x)=\chi(x)\sum_{j\in\mathbb{N},\,k\in\mathbb{Z}}b_{j,k}(\theta^{n}_{j,k,1}+i\theta_{j,k,2}^{n})\alpha_{j}(t)e_{k}(x),\quad t\in[0,1).$
    Here $\chi\in C^{\infty}(\mathbb{T})\setminus\{0\}$ is an arbitrary non-zero function333[7] assumed $\chi$ to be positive in an interval. But the proof still works out as long as $\chi$ is not identically zero., $b_{j,k}\geq 0$ are deterministic constants tending to 0 sufficiently fast (see (1.2)), and $\theta_{j,k,l}^{n}$ are independent real-valued random variables. Moreover, $\theta_{j,k,l}^{n}$ admits a probability density function $\rho_{j,k,l}$ supported by the interval [1,1][-1,1], which is $C^{1}$ and satisfies $\rho_{j,k,l}(0)>0$.
In view of the boundedness condition (1.2), $\eta_{n}$ are i.i.d. $L^{2}(0,1;H^{1+\sigma}(\mathbb{T}))$-valued random variables. Moreover, their common law $\mathscr{D}(\eta)\in\mathcal{P}(L^{2}(0,1;H^{1+\sigma}(\mathbb{T})))$ has compact support $E:=\operatorname{supp}(\mathscr{D}(\eta_{n}))$, due to a diagonal argument (or Tikhonov’s theorem). We emphasize that $E$ is equipped with the topology of $L^{2}(0,1;H^{1+\sigma})$, rather than the weaker one of $L^{2}(0,1;H^{1})$.
@@
$S\colon H^{1}\times L^{2}(0,1;H^{1})\to H^{1},\quad S(u_{0},\zeta)=u(1).$
@@
Hypothesis $(\mathbf{C})$ has been demonstrated in [7, Section 5.3] via control theory, provided we equip $X$ with an equivalent norm (see [7, Lemma 4.10]). We mention that the non-degeneracy condition (1.3) and the $L^{2}(0,1;H^{1+\sigma})$-compactness of $E$ are crucial to this point.
Meanwhile, a weaker version of $(\mathbf{I})$ is also derived in this prior work by global stability. Indeed, [7, Section 5.3] shows that (2.2) holds when $y=0$; that is, for any $\varepsilon_{0}>0$, one can find $N_{0}\in\mathbb{N}$ and $p_{0}>0$ such that
@@
and the sequence $\mathcal{A}_{n}(\{0\})$ is increasing (since $S(0,0)=0$ and $0\in E$). By compactness we see that $X\subset B_{X}(\mathcal{A}_{N_{1}}(\{0\}),\varepsilon_{1})$ for some $N_{1}\in\mathbb{N}$. In other words, for any $y\in X$, we can find $\zeta_{0},\dots,\zeta_{N_{1}-1}\in E$ such that
@@
Thus, since $\eta_{n}$ are i.i.d., we find that for any $z\in B_{X}(0,\delta)$,
@@
Since $\operatorname{supp}(\mathscr{D}(\eta_{n}))=E\ni\zeta\mapsto\mathbb{P}(\eta_{n}\in B_{E}(\zeta,\delta))$ is lower-semicontinuous and pointwise positive, and $E$ is compact, we conclude that $p_{1}$ is a positive constant. Finally, for any $\varepsilon>0$, let us set $\varepsilon_{1}=\varepsilon/2$, which determines the constant $\delta$, and then set $\varepsilon_{0}=\delta$. Then for $N:=N_{0}+N_{1}$ and any $x,y\in X$, by Markov property, we have
$P_{N}(x,B_{X}(y,\varepsilon))\geq P_{N_{0}}(x,B_{X}(0,\delta))\inf_{z\in B_{X}(0,\delta)}P_{N_{1}}(z,B_{X}(y,\varepsilon))\geq p_{0}p_{1}>0.$
@@ A $\Gamma$-type domain is a subdomain of $D$ in the form
    $\eta(t,x)=\eta_{n}(t-n,x)\quad\text{for }t\in[n,n+1),\ n\in\mathbb{N}_{0},$ (2.7)
@@ 2.3.2. Navier–Stokes system
Finally, we investigate the 2D Navier–Stokes system with bounded kick force. In particular, our result extends [21, Theorem A] by allowing the noise to degenerate in high frequencies. The system in consideration is studied in, e.g., [26, 21, 22], which reads
@@
Here $D$ is a smooth bounded domain in $\mathbb{R}^{2}$, with boundary $\partial D$ and outer normal vector $n(x)$, the constant $\nu>0$ is the viscosity, $u=(u^{1},u^{2})$ is the velocity field, and $p$ is the pressure. The phase space for (2.9) is specified as
@@
    $\eta(t,x)=\sum_{n=1}^{\infty}\delta(t-n)\eta_{n-1}(x),\quad\eta_{n}(x)=\sum_{k=1}^{\infty}b_{k}\theta_{k}^{n}\varphi_{k}(x),$ (2.10)
@@
Finally, we mention a similar application to Navier–Stokes system perturbed by degenerate space-time localized bounded noise, which has been studied in [35] for exponential mixing. In this case the noise is akin to our setting for Schrödinger and wave equations. For the sake of simplicity, we omit the detailed description of this example, and claim that the LDP holds for the empirical distributions, as both $(\mathbf{I})$ and $(\mathbf{C}$) has essentially been verified in that work.
@@ 3.1. Some reductions
Reduction 1. Kifer’s criterion. In order to investigate the uniform LDP, let us consider the directed set [^6] A directed set $(\Theta,\prec)$ is a non-empty set $\Theta$ equipped with a pre-order $\prec$ (i.e. a reflexive and transitive binary relation), such that any finite subset admits an upper bound. A net $\{x_{\theta}:\theta\in\Theta\}$ in a topological space $X$ refers to elements labeled by this directed set. The net $x_{\theta}$ converges to $x$, if for any neighborhood $U$ of $x$, there exists $\theta_{0}\in\Theta$ such that $x_{\theta}\in U$ whenever $\theta_{0}\prec\theta$. $\Theta:=\mathbb{N}\times\mathcal{P}(X)$, endowed with the pre-order $\prec$ defined by
@@
which is lower semicontinuous and convex. Moreover, $\Lambda$ can be reconstructed by the dual relation:
@@
Let $\Lambda\colon C(X)\to\mathbb{R}$ and $I\colon\mathcal{P}(X)\to[0,\infty]$ be defined as above. Given $V\in C(X)$, if the supremum on the right-hand side of (3.3) is attained at $\sigma_{V}\in\mathcal{P}(X)$, then $\sigma_{V}$ is called an equilibrium state for $V$.
@@
Assume the hypotheses $(\mathbf{I})$ and $(\mathbf{C})$ are satisfied, and $V\in L(X)$. Then there exists $\lambda_{V}>0$, $h_{V}\in C(X)$ and $\mu_{V}\in\mathcal{P}(X)$, such that for any $f\in C(X)$ and $\nu\in\mathcal{M}_{+}(X)$,
@@ 3.2. Lipschitz estimates for $Q_{n}^{V}$
Since $X$ is compact, so is $\mathcal{P}(X)$. Then Schauder’s fixed point theorem implies that there exists a fixed point $\mu\in\mathcal{P}(X)$. In other words, let $\lambda_{2}=(Q^{V*}\mu)(X)>0$, we get (3.18).
@@ 4. Further discussion
Acknowledgments   The authors thank the anonymous referee for insightful comments and suggestions. We are also grateful to Ziyu Liu for valuable discussions and advices during the preparation of the paper. Yuxuan Chen would like to appreciate his PhD advisor Zhifei Zhang. Shengquan Xiang is partially supported by NSFC 12301562.
@@ B.  Asymptotics of generalized Markov semigroups
Then there exists $\lambda>0$, $h\in C(X)$ and $\mu\in\mathcal{P}(X)$, such that for any $f\in C(X)$ and $\nu\in\mathcal{M}_{+}(X)$,